OpenStudy (anonymous):

how to I solve a question like: m over pi. = 7.3??? Thanks P.S will give medal for best answer :)

hartnn (hartnn):

\[\frac{m}{\pi}=7.3\] m=7.3*pi do u know the value of pi??

OpenStudy (anonymous):

the value is 3.14159265359

hartnn (hartnn):

yes,plug in this value in 7.3 * pi and find m

OpenStudy (anonymous):

THANKS soooo much :)

hartnn (hartnn):

your welcome :)